### v1

# Defect of irreducible plane curves with simple singularities

Abstract
In this note we focus on the defect of singular plane curve that was recently introduced by Dimca. Roughly speaking, the defect of a reduced plane curve measures the discrepancy from the property of being a free curve. We find some lower-bound on the defect for certain classes of irreducible plane curves admitting nodes, ordinary cusps and ordinary triple points. The main result of the note tells us that reduced simply singular plane curves with sufficiently high Arnold exponents are never free.
Keywords defect; simply singular plane curves
Mathematics Subject Classification (2020) 14N25, 14H50, 32S25
To Professor Arkadiusz Płoski, in memoriam.
In this note we study the notion of the defect for some classes of irreducible plane curves. This notion has been introduced by Dimca in [4] and we have many extremely interesting questions revolving around this notion that we may want to study. In order to present a sample of possible questions, we need a solid preparation that is based on [3].
Let $S:=\mathbb{C}[x,y,z]=\bigoplus_{k}S_{k}$ be the graded polynomial ring and let $f\in S$ be a homogeneous polynomial of degree $d$. Consider a reduced curve $C\,:\,f=0$ in $\mathbb{P}^{2}_{\mathbb{C}}$ defined by $f$. We denote by $\partial_{x},\partial_{y},\partial_{z}$ the partial derivatives and we define ${\rm Der}(S)=\{\partial:=a\cdot\partial_{x}+b\cdot\partial_{y}+c\cdot\partial_{z},\,\,a,b,c\in S\}$ which is the free $S$-module of $\mathbb{C}$-linear derivations of the ring $S$. Now for a reduced curve $C\,:f=0$ with $f\in S_{d}$ being homogeneous, we define
${\rm D}(f)=\{\partial\in{\rm Der}(S)\,:\,\partial(f)\in\langle f\rangle\}.$
It means that ${\rm D}(f)$ is the graded $S$-module of derivations preserving the ideal $\langle f\rangle$. One can show that for a reduced curve $C\,:f=0$ in $\mathbb{P}^{2}_{\mathbb{C}}$ we have the following decomposition:
${\rm D}(f)={\rm D}_{0}(f)\oplus S\cdot\delta_{E},$
where $\delta_{E}=x\partial_{x}+y\partial_{y}+z\partial_{z}$ is the Euler derivation, and
${\rm D}_{0}(f)=\{\partial\in{\rm Der}(S)\,:\,\partial\,f=0\},$
i.e., the set of all $\mathbb{C}$-linear derivations of $S$ killing the polynomial $f$. It is classically known that ${\rm D}_{0}(f)$ can be identified with the $S$-module of all non-trivial Jacobian relations for the partials of $f$, namely
${\rm AR}(f)=\{(a,b,c)\in S^{3}\,:\,a\cdot\partial_{x}\,f+b\cdot\partial_{y}\,f+c\cdot\partial_{z}\,f=0\}.$
We have some important numerical invariants that one can associated with a curve $C\,:f=0$ in $\mathbb{P}^{2}_{\mathbb{C}}$, one of them is the minimal degree among derivations killing $f$, i.e.,
${\rm mdr}(f)={\rm min}\{r\in\mathbb{N}\,:\,{\rm D}_{0}(f)_{r}\neq 0\}={\rm min}\{r\in\mathbb{N}\,:\,{\rm AR}(f)_{r}\neq 0\}.$
Sometimes we will write ${\rm mdr}(C)$ for a given curve $C\subset\mathbb{P}^{2}_{\mathbb{C}}$.
For a homogeneous polynomial $g\in S$ of degree $d$ we define its Jacobian ideal $J_{g}:=\langle\partial_{x}\,g,\partial_{y}\,g,\partial_{z}\,g\rangle$, and we define by $I_{g}$ the saturation of $J_{g}$ with respect to the irrelevant ideal $\mathfrak{m}=\langle x,y,z\rangle$. The Jacobian module of $g$ is defined as
$N(g)=I_{g}/J_{g}.$
The Jacobian module provides important information about the curve that is defined by $f\in S$. In order to show its strength, let us introduce the following definition.
Definition 1.
A reduced curve $C\,:f=0$ in $\mathbb{P}^{2}_{\mathbb{C}}$ defined by a homogeneous polynomial $f\in S$ is free if ${\rm D}(f)$, or just ${\rm D}_{0}(f)$, is a free graded $S$-module.
It turns out that the freeness of $C\,:f=0$ boils down to the condition $N(f)=0$, i.e., the Jacobian ideal is saturated. We set
$n(f)_{j}={\rm dim}\,N(f)_{j},$
and for a reduced curve $C\,:f=0$ in $\mathbb{P}^{2}_{\mathbb{C}}$ given by $f\in S$ we define the following invariant
$\nu(C)={\rm max}\{n(g)_{j}\}_{j}.$
The invariant $\nu(C)$ is called the defect, or the defect from the freeness property. It is very difficult to compute the defect of a given curve $C$ by using the above definition. However, Dimca showed the following crucial result. Recall that for a reduced plane curve $C\,:f=0$ we denote by $\tau(C)$ its total Tjurina number, i.e.,
$\tau(C)=\sum_{p\in{\rm Sing}(C)}\tau_{p}(C),$
where ${\rm Sing}(C)$ denotes the set of all singular points of $C$ and $\tau_{p}(C)$ the local Tjurina number of $p$.
Theorem 2.
([4, Theorem 1.2]) If $C\,:\,f=0$ is a reduced plane curve of degree $d$ and $r={\rm mdr}(f)$.
  • If $r<(d-1)/2$, then $\nu(C)=(d-1)^{2}-r(d-1-r)-\tau(C)$.
  • If $r\geqslant(d-2)/2$, then
    $\nu(C)=\bigg{\lceil}\frac{3}{4}(d-1)^{2}\bigg{\rceil}-\tau(C).$
There are many interesting and difficult open problems regarding the notion of the defect and here we would like to recall two the most important conjectures. The first one can be considered as a vast generalization of the Terao’s freeness conjecture and it is devoted to line arrangements.
Conjecture 3.
For a line arrangement $\mathcal{L}\subset\mathbb{P}^{2}_{\mathbb{C}}$ the defect $\nu(\mathcal{L})$ is determined by the intersection lattice of $\mathcal{L}$. More precisely, if $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$ are two line arrangements that have isomorphic intersection lattices, then $\nu(\mathcal{L}_{1})=\nu(\mathcal{L}_{2})$.
This conjecture seems to be extremely difficult and for more details about it we refer the reader to an excellent recent survey by Dimca [5]. In the case of our note, we focus on the case of irreducible plane curves, and in order to present the main motivation for our research we need two additional definitions.
Definition 4.
A plane rational cuspidal curve is a rational curve $C:f=0$ in $\mathbb{P}^{2}_{\mathbb{C}}$ having only unibranch singularities.
It is also necessary to introduce another important class of curves.
Definition 5.
A reduced curve $C\,:f=0$ in $\mathbb{P}^{2}_{\mathbb{C}}$ defined by a homogeneous polynomial $f\in S$ is nearly free if $\nu(C)=1$.
In the light of the above definitions, we have the following truly surprising conjecture.
Conjecture 6.
Any rational cuspidal curve $C$ is either free or nearly free.
In the present note, strongly motivated by the above conjecture, we want to continue the idea of studying the defect for some natural classes of irreducible plane curves, since, apart from the above conjecture, we do not have any general prediction or results devoted to such curves.
Our first result is devoted to nodal curves.
Definition 7.
We say that an irreducible and reduced curve of $C_{d}\subset\mathbb{P}^{2}_{\mathbb{C}}$ of degree $d$ is nodal if admits only $n_{2}\leqslant\frac{(d-1)(d-2)}{2}$ nodes as singularities.
Theorem A.
Let $C_{d}$ be a nodal plane curve of degree $d\geqslant 4$. Then
$\nu(C_{d})\geqslant\frac{1}{4}(d^{2}-1).$
In particular, the defect for nodal curves can be arbitrarily large.
The next result is devoted to irreducible and reduced plane curves of genus zero admitting only nodes and ordinary triple points as singularities.
Theorem B.
Let $K_{3k}$ be an irreducible and reduced plane curve of degree $d=3k$ with $k\geqslant 3$ of genus zero that admits exactly $2k$ ordinary triple points. Then
$\nu(K_{3k})\geqslant\frac{1}{4}(9k+1)(k-1).$
Finally, we focus on certain cuspidal curves that very constructed by Ivinskis [9].
Theorem C.
Let $C_{6k}$ be an irreducible and reduced plane curve of degree $d=6k$ with $k\geqslant 1$ that admits exactly $9k^{2}$ ordinary cusps. Then
$\nu(C_{6k})=9k^{2}-9k+1=g(C_{6k}).$
In particular, $C_{6}$ is nearly free.
Our results show that rational cuspidal plane curves are very special and it allows us to justify a bit heuristic phenomenon that it is very difficult to construct irreducible free or nearly free curves.
Before we present the proofs, we need to recall major tools that we are going to use here. We start with the following crucial result [6, Theorem 2.1].
Theorem 8 (Dimca-Sernesi).
Let $C\,:\,f=0$ be a reduced curve of degree $d$ in $\mathbb{P}^{2}_{\mathbb{C}}$ having only quasi-homogeneous singularities. Then
${\rm mdr}(f)\geqslant\alpha_{C}\cdot d-2,$
where $\alpha_{C}$ denotes the Arnold exponent of $C$.
The Arnold exponent of a given reduced curve $C\subset\mathbb{P}^{2}_{\mathbb{C}}$ is defined as the minimum over all log canonical thresholds ${\rm lct}_{p}(C)$ for $p\in{\rm Sing}(C)$. In the case when our singularities are just ordinary, we have the following result [2, Theorem 1.3].
Theorem 9.
Let $C$ be a reduced curve in $\mathbb{C}^{2}$ which has degree $m$. Then ${\rm lct}_{p}(C)\geqslant\frac{2}{m}$, and the equality holds if and only if $C$ is a union of $m$ lines passing through 00.
By the above result, if $p=(0,0)\in\mathbb{C}^{2}$ is an ordinary singularity of multiplicity $r$ of $C$, then
${\rm lct}_{p}(C)=\frac{2}{r}.$ (1)
Furthermore, if $q\in C$ is an ordinary cusp having the local normal form $y^{2}+x^{3}=0$, then by [2, Example 1.5] we have
${\rm lct}_{q}(C)=\frac{5}{6}.$ (2)
Now we are ready to present our proof of Theorem A.
Proof.
Let us recall that by a result due to Severi [11] there exist irreducible nodal curves $C_{d}$ of degree $d\geqslant 3$ having exactly $n_{2}\leqslant(d-1)(d-2)/2$ nodes. Since all singular points $p\in{\rm Sing}(C)$ are nodes, then ${\rm lct}_{p}(C_{d})=1$ and the Arnold exponent of $C_{d}$ is equal to
$\alpha_{C_{d}}=1.$
Then by Theorem 8 we have
${\rm mdr}(C_{d})\geqslant d-2.$
By the assumption $d\geqslant 4$, and the following inequality holds
$d-2>\frac{d-2}{2},$
which means that the defect of $C_{d}$ is can be computed as
$\nu(C)=\bigg{\lceil}\frac{3}{4}(d-1)^{2}\bigg{\rceil}-\tau(C).$
Observe that $\tau(C_{d})\leqslant\frac{(d-1)(d-2)}{2}$, and we finally get
$\nu(C_{d})\geqslant\frac{3}{4}(d-1)^{2}-\frac{(d-1)(d-2)}{2}=\frac{1}{4}(d^{2}-1),$
which completes the proof. ∎
Now we pass to our proof of Theorem B.
Proof.
The existence of such irreducible curves $K_{3k}$ of genus zero with $n_{3}=2k$ triple points is granted by [8, 3.4 Theorem]. The condition that $K_{3k}$ has genus zero means that the curve has exactly
$n_{2}=\frac{(3k-1)(3k-2)}{2}-3\cdot 2k=\frac{9k^{2}-21k+2}{2}$
nodes as singularities. Our curve $K_{3k}$ admits only nodes and ordinary triple points as singularities which implies that
$\alpha_{K_{3k}}={\rm min}\bigg{\{}1,\frac{2}{3}\bigg{\}}=\frac{2}{3},$
and then by Theorem 8
${\rm mdr}(K_{3k})\geqslant\frac{2}{3}\cdot 3k-2=2k-2.$
Since $k\geqslant 3$, we have
$2k-2>\frac{3k-2}{2},$
so the defect of $K_{3k}$ can be bound from below as
$\nu(K_{3k})=\bigg{\lceil}\frac{3}{4}(3k-1)^{2}\bigg{\rceil}-4\cdot 2k-\frac{9k^{2}-21k+2}{2}\geqslant\frac{1}{4}(9k+1)(k-1),$
which completes the proof. ∎
Finally, we present our proof of Theorem C.
Proof.
Let us start with the geometric construction of $C_{6k}$ with $k\geqslant 1$. In his Diplomarbeit, Ivinskis shows that there exists an irreducible and reduced curve $C_{6k}$ of degree $6k$ with $k\geqslant 1$ having exactly $9k^{2}$ ordinary cusps [9, Lemma 4.1.7]. This curve is constructed using the Kummer cover $\kappa:\mathbb{P}^{2}\ni(x,y,z)\mapsto(x^{n},y^{n},z^{n})\in\mathbb{P}^{2}$ applied to an irreducible and reduced sextic with exactly $9$ ordinary cusps. Recall that such an irreducible sextic is the dual curve to a smooth elliptic curve $E$, and the ordinary cusps correspond to the $9$ inflection points of $E$. Since our curve $C_{6k}$ admits only ordinary cusps as singularities, then
$\alpha_{C_{k}}=\frac{5}{6}$
and by Theorem 8 we have
${\rm mdr}(C_{6k})\geqslant\frac{5}{6}\cdot 6k-2=5k-2.$
Since for $k\geqslant 1$ one has
$5k-2>\frac{6k-2}{2}=3k-1,$
and $\tau(C_{6k})=2\cdot 9k^{2}=18k^{2},$ the defect of $C_{6k}$ is equal to
$\nu(C_{6k})=\bigg{\lceil}\frac{3}{4}(6k-1)^{2}\bigg{\rceil}-18k^{2}.$
Observe that
$\bigg{\lceil}\frac{3}{4}(6k-1)^{2}\bigg{\rceil}=\bigg{\lceil}27k^{2}-9k+\frac{3}{4}\bigg{\rceil}=27k^{2}-9k+1,$
and then
$\nu(C_{6k})=27k^{2}-9k+1-18k^{2}=9k^{2}-9k+1.$
In particular, for $k=1$ our curve $C_{6k}$ is an irreducible sextic with $9$ ordinary cusps with $\nu(C_{6})=1$, so $C_{6}$ is nearly free. ∎
Remark 10.
Our curves $C_{6k}$ considered above are obviously not rational since
$g(C_{6k})=\frac{(6k-1)(6k-2)}{2}-9k^{2}=9k^{2}-9k+1\geqslant 1.$
Moreover, it shows that $g(C_{6k})=\nu(C_{6k})$, which is extremely surprising.
Let us now present the main result of the note. This result is devoted to reduced simply singular plane curves, i.e., reduced plane curves with only ${\rm ADE}$ singularities.
Theorem D (Non-freeness criterion).
Let $C\subset\mathbb{P}^{2}_{\mathbb{C}}$ be a reduced plane curve of even degree $d=2m\geqslant 4$ admitting only ${\rm ADE}$ singularities. Assume furthermore that the Arnold exponent of $C$ satisfies $\alpha_{C}\geqslant\frac{1}{2}+\frac{1}{m}$. Then
$\nu(C)\geqslant 1.$
In particular, $C$ is never free.
Proof.
The condition that $\alpha_{C}\geqslant\frac{1}{2}+\frac{1}{m}$ ensures us that the defect of $C$ can be computed via the formula
$\nu(C)=\bigg{\lceil}\frac{3}{4}(2m-1)^{2}\bigg{\rceil}-\tau(C).$
Observe that
$\bigg{\lceil}\frac{3}{4}(2m-1)^{2}\bigg{\rceil}=3m^{2}-3m+1,$
so the last thing that we need to estimate is $\tau(C)$. Since ${\rm mdr}(C)\geqslant m$, which follows from the fact that $\alpha_{C}\geqslant\frac{1}{2}+\frac{1}{m}$, then by a result due to Du Plessis and Wall in [7] we see that
$\tau(C)\leqslant\tau_{{\rm max}}(2m,r):=(2m-1)(2m-r-1)+r^{2}-\binom{2r-2m+2}{2},$
where $r:={\rm mdr}(C)$. Since the function $\tau_{{\rm max}}(2m,r)$ is strictly decreasing as a function with respect to $r$ on the interval $[m,2m-1]$, then
$\tau(C)\leqslant\tau_{{\rm max}}(2m,m)=3m^{2}-3m,$
so we finally get
$\nu(C)=3m^{2}-3m+1-\tau(C)\geqslant 3m^{2}-3m+1-\tau_{{\rm max}}(2m,m)=1,$
which completes the proof. ∎
Finally, we present the following example to show that our main result is sharp in the strict sense.
Example 11.
Fix an even integer $m\in\mathbb{Z}_{\geqslant 4}$ and consider the curve $C_{2m}=\{C_{1},C_{2},C_{3},C_{4}\}\subset\mathbb{P}^{2}_{\mathbb{C}}$, where
$\begin{array}[]{l}C_{1}:\quad x^{m/2}+y^{m/2}+z^{m/2}=0,\\ C_{2}:\,-x^{m/2}+y^{m/2}+z^{m/2}=0,\\ C_{3}:\quad x^{m/2}-y^{m/2}+z^{m/2}=0,\\ C_{4}:\quad x^{m/2}+y^{m/2}-z^{m/2}=0.\end{array}$
Our curve $C_{2m}$ is of degree $d=2m$ and it has $3m$ singularities of type $A_{m-1}$, see [10, Lemma 7.5]. In particular, for $m=4$ we obtain the arrangement of $4$ conics that admits exactly $12$ singularities of type $A_{3}$ – it is well-known thaty this arrangement is unique up to the projective equivalence. Since $C_{2m}$ admits only singularities of type $A_{m-1}$, then the Arnold exponent of $C_{2m}$ is equal to
$\alpha_{C_{2m}}=\frac{1}{2}+\frac{1}{m},$
and this follows from the fact that for each $p\in{\rm Sing}(C_{2m})$ one has ${\rm lct}_{p}=\frac{1}{2}+\frac{1}{m}$.
It means that by our main result we have
$\nu(C_{2m})\geqslant 1.$
In fact, based on [1, Theorem 3.12], our curve $C_{2m}$ is nearly free, i.e., $\nu(C_{2m})=1$, which explains why our result is sharp.
Acknowledgement
I would like to thank Emilia Mezzetti for useful explanations regarding the content of [8].
Piotr Pokora is supported by the National Science Centre (Poland) Sonata Bis Grant 2023/50/E/ST1/00025. For the purpose of Open Access, the author has applied a CC-BY public copyright licence to any Author Accepted Manuscript (AAM) version arising from this submission.